The sum of two numbers is 60.One number is 16 more than twice the other number.
Answer by hovuquocan1997(83)

let's the second number be x, the first number is 16 more than twice the second number, so it will be 2x+16
The sum of two number is 60, means x + 2x + 16 = 60
3x = 44
x = 44/3
So one number is 44/3 and the other one is 2*44/3 + 16 = 136/3
TA-DAH
:D